Find "X" and show how please
2/5 X = 4/3
X=?

Answer by alittletooconfused(7)

2/5 X = 4/3
X=?
Solution:
Multiply both sides by 5
5(2/5 X = 4/3)5
Result will be,
2X = 20/3
We need to cancel 2 so that only X will be left. In that case, we divide both sides by 2
2X/2 = 20/3 ÷ 2
Result will be,
X = 10/3